A mutal fund with 1 million shares has 800 million of assets and 20 million of liabilities. What is its NAV? If fund is traded at $55 per share what is it premium or discount in dollars and as percentage? What type of fund it must be if its price is (substantially) different from NAV?
Answer #1

NAV=(Assets-Liabilities)/Number of Shares=(800-20)/1=78

Discount in dollars=78-55=-23

Discount in %=23/78=29.49%

As price is substantially different from NAV, this is a closed end fund